# Robust Containment Queries over Collections of Trimmed NURBS Surfaces via Generalized Winding Numbers

### v2

@@ 3.1. Algorithm Overview
Our algorithm is derived from Stokes’ theorem, which reduces the dimension of Equation 5 to an integral over the boundary of each surface. For notational simplicity, we assume that the surface $S$ is translated so that the query point $q$ is at the origin.
In summary, for each translated surface $S$ of the model, we evaluate the GWN at query point $q=0$ by evaluating
@@
The first term is our Stokes’-derived boundary formulation, which we evaluate using an error-controlled adaptive Gaussian quadrature scheme. This evaluation is accelerated across batches of queries through a memoization strategy which caches and reuses geometric data for each surface, such as surface points and tangent vectors at quadrature nodes (as detailed in Section 3.4). The second term is a correction term that accounts for any jump discontinuities in the GWN field over the surface.
@@
  • Far-field: For points sufficiently far from the surface, i.e. outside a bounding box encompassing the surface, the number of intersections is known to be zero. As such, we evaluate the GWN directly with our un-adjusted boundary formulation.
@@
The specific form of Equation 7 assumes that correction term is identified from intersections with the $z$-axis. However, the intersection test can be performed with an arbitrary line by subsequently rotating the surface so that the line is parallel to the $z$-axis. As intersections with lines that are near-tangent to the surface are the most difficult to identify and resolve, we heuristically select a line with a direction equal to an average surface normal, which we compute for each surface as preprocessing. In cases where the symmetry of the surface results in an average normal with near-zero magnitude, we instead select the line randomly.
@@ 3.2. Reformulation with Stokes’ Theorem
Evaluating such an integral via numerical quadrature along the lower-dimensional space naturally improves the efficiency of the total method. This is a well-established advantage of Stokes’ theorem, and is the basis of a method proposed by Gunderman et al. (2021) for evaluating arbitrary integrals on trimmed NURBS surfaces. Here, we make the approach even more practical by evaluating the “antiderivative” $F$ exactly for the relevant $\nabla\times F$, further reducing the quadrature nodes in a scheme of fixed order $n$ from $O(n^{2})$ to $O(n)$.
Our Stokes’-based reformulation has the added benefit of fundamentally improving the stability of the ensuing quadrature scheme, as query points which are near to the surface but not near the boundary are not subject to the same near-singular behavior in the corresponding integrand. Furthermore, we capitalize on this boundary reformulation through our use of a geometrically adaptive quadrature method, the computational burden of which is dramatically reduced by considering subdivision only of 3D space curves.
@@
(9) $\nabla\times F=\frac{1}{4\pi}\cdot\frac{x}{\lVert x\rVert^{3}}.$
Recall that we have made the assumption that the query point $q$ is located at the origin, and so this integrand is continuous on the surface whenever $x\neq 0$, i.e. the query point is not coincident with the surface. (We treat this case explicitly in Section 3.3).
Given this, there are many distinct vector fields that satisfy Equation 9, and we consider one such field in Cartesian coordinates:
(10) $F(x,y,z)=\frac{1}{4\pi}\left\langle\frac{yz}{(x^{2}+y^{2})\,\lVert\vec{x}\rVert},\frac{-xz}{(x^{2}+y^{2})\,\lVert\vec{x}\rVert},0\right\rangle.$
If the theoretical conditions of Stokes’ theorem can be met, then we can evaluate the GWN for an arbitrary surface with the following 1D line integral,
@@
a reformulation of the problem statement that comes with a number of computational advantages.
However, the formulation of $F$ in Equation 11 introduces theoretical difficulties. For the conditions of Stokes’ theorem to be met, both the original vector field and its antiderivative must be continuous—not just on its domain of integration (the boundary of the surface) but also on the original domain of integration (the surface itself). Notably, this antiderivative is discontinuous whenever $x^{2}+y^{2}=0$ on the surface, i.e. whenever the surface intersects a line extending in both $z$ directions from the query point. In such cases, we cannot mathematically justify the direct use of Stokes’ theorem, even if the resulting line integral is itself continuous.
We note that particular antiderivative $F$ in Equation 10, characterized by a line of discontinuities on the $z$-axis, was chosen arbitrarily. Indeed, one can analogously define a family of antiderivatives $F$, each characterized by a different line of discontinuities.
@@
In Figure 5, we see examples of different configurations of the surface and the line. As shown in Figure 5(b), it is practical to consider different choices for $F$ not as explicit vector fields, but rather as a rotation of the NURBS patch via its control points so that the desired line of discontinuities is aligned to the $z$-axis. In this way, applying Stokes’ theorem requires choosing an antiderivative $F$ whose corresponding line of discontinuities does not intersect the surface, and rotating the surface accordingly. In practice, one cannot always identify such a line for query points which are nearer to the surface, nor does one necessarily exist (see Figure 5(c)).
@@ 3.2.1. Far-field: Direct Application of Stokes’ Theorem
Other convex shapes containing the shifted surface can also be used to guarantee a non-intersecting line $L$, again followed by the appropriate rotation of the surface. For example, our implementation also checks a precomputed oriented bounding box for each NURBS surface, allowing a greater number of points to be classified as far-field. In any case, these are the most commonly evaluated forms of the integral, as an arbitrary query point will be considered far from all but a few surfaces in a typical CAD model.
@@ 3.2.2. Near-field: Analytically Adjusted Stokes’ Theorem
If the surface bounding box does contain the origin (i.e. the query point is inside the surface bounding box), we can no longer guarantee that there exists a line $L$ which does not intersect the surface. Instead, we must explicitly test for intersection between some $L$ and the surface. We consider cases to be “near-field” when all intersections between $L$ and the surface occur away from the surface edges and other degeneracies. When these conditions are not met, the query point is handled as an “edge-case”.
@@
Naturally, distinguishing the near-field and edge cases requires first identifying the location of all intersections between the surface and the line in the parameter space of the surface. As noted in Section 2, there are many numerical ray-casting algorithms one could apply to find intersections between this line and the surface, and in principle, any can be used for the purposes of evaluating the GWN. We describe our own geometric subdivision approach in Section 5.1.1, which in contrast to many algebraic alternatives, naturally takes advantage of the broader GWN algorithm’s high tolerance for imprecision in the intersection routine.
@@
Because any intersection between $L$ and the surface violates the conditions of Stokes’ theorem, direct application of the boundary formulation is theoretically unfounded and, in general, evaluates to an incorrect GWN. Our proposed correction term in Equation 7 permits evaluation of the GWN using the same boundary integral formulation as Equation 11, provided that for each intersection (if any), the value $0.5$ is added or subtracted according to the orientation of the surface at the point of intersection to compensate for the jump discontinuity present in the GWN field along the surface.
@@
This is illustrated in Figure 6, which shows that the difference between the ground truth GWN and the specific boundary formulation in Equation 11 is always half-integer valued, and in turn how the ground truth GWN can always be recovered from the boundary formulation by identifying the proper half-integer value.
Figure 6. The difference between the 3D GWN for a curved surface (left, shown at two planar slices) and that of a particular boundary term (middle) is always half-integer valued, and partitioned by the geometry (right).
In some sense, the existence of this analytic fix for the discontinuity in the integral is a natural consequence of the double-layer potential which defines the GWN scalar field (Barill et al., 2018; Bang et al., 2023). However, the use of this specific correction term is made more intuitive by means of analogy in two spatial dimensions (with a direct proof provided in Appendix A).
@@
As summarized in Section 2.3 and demonstrated in Figure 7, the GWN for an arbitrary curve in 2D (left) can always be computed as the difference between an integer winding number for the curve closed by a straight line (right) and a generalized winding number defined by that straight line closure (middle). Crucially, the GWN of the closure is uniquely defined by the two boundary points of the original curve. From the lens of our 3D GWN framework, the GWN of the closure can be thought of as a “boundary term” which must be “corrected” by the integer winding number of the closed shape.
@@
To make the comparison explicit, we see that in both 2D and 3D, the value of the GWN is the sum of a fractional component which does not depend on any interior geometry, and a discrete component which does. Although the fractional component in 2D reduces to a subtended angle, while the 3D equivalent is not the GWN of any particular shape, both discrete components can be thought of as determined by the result of a line intersection test. Most importantly, points which are known to be far from the shape do not require any correction at all. Returning our focus to 3D, we emphasize that although our line-surface intersection test does come at additional cost, it is necessary only for the surfaces nearest to the query point.
Figure 7. The difference between the 2D GWN for a curve (left) and that of a particular boundary term (i.e. its linear closure) (middle) is always integer valued, and partitioned by the geometry (right).
@@ 3.2.3. Edge Cases: Disk Extraction via Trimming Curves
We now consider the remaining edge cases for which the query point is inside the surface bounding box, but the correction term in Equation 7 cannot be applied. As shown in Figure 8, this occurs when the line of discontinuities intersects the surface at a boundary, at a cusp, or is tangent to the surface. We treat all edge cases by either finding a new line (in the case of tangencies) or by removing from the surface a disk of radius $r$ in parameter space around the point of intersection. This guarantees that the line does not intersect the surface at the otherwise problematic intersection point, and permits evaluation via Equation 7. The extracted disk is then fed back into the original algorithm for re-processing. Since the produced disk has a much smaller bounding box than the original surface, the computational cost of the re-processing is significantly reduced.
@@
The only edge case which is not well-handled by this strategy is the case of a line that is tangent (or near-tangent) to the surface, as the surface remaining after the disk is removed may still intersect the line. In these cases, we instead rotate the surface (i.e. consider a different line) and feed the rotated surface into the original algorithm (See Figure 8(c)). While this is the most computationally expensive type of edge case, it is also the least common, in large part due to our careful selection of the cast ray to be an approximate surface normal (See Section 5.1.1).
@@
3.3. GWN for Coincident Points
@@ 3.4. Accelerated Adaptive Quadrature
In our implementation, we utilize a recursive subdivision approach, applying a fixed-order Gauss-Legendre quadrature rule at each level of recursion. We stop the recursion when the difference between the evaluated integral on the entire curve and the sum of the integrals on the two halves is less than a user-defined quadrature tolerance $\epsilon_{q}$. Thus, in most cases, $\epsilon_{q}$ determines how many digits of precision are achieved in the computed GWN field, although it is possible for more lenient values of $\epsilon_{q}$ that the two terms will agree at an incorrect value. Using a Gaussian rule at each recursion level is particularly useful as they are capable of maintaining high precision even without specific structure in the integrand. Furthermore, our recursive subdivision strategy is specifically chosen because it effectively concentrates quadrature nodes near unstable regions of the integrand, i.e. when the query point is close to the 3D space curve which defines the domain of integration. Elsewhere, where the integrand is stable, we maintain the typically high precision of Gaussian quadrature without additional computational expense.

### v3

@@ 3.1. Algorithm Overview
Our algorithm is based on a direct application of Stokes’ theorem, which reduces the dimension of Equation 4 to an integral over the boundary of each surface. For notational simplicity, we assume that the surface $S$ is translated so that the query point $q$ is at the origin.
For each translated surface $S$ of the model, we evaluate the GWN at query point $q=0$ by evaluating
@@
where $q$ is on the positive side of $S$ if for intersection $(0,0,z_{0})$, $z_{0}>0$.
The first term is the Stokes’-derived boundary formulation, which we evaluate using an error-controlled adaptive Gaussian quadrature scheme. This evaluation is accelerated across batches of queries through a memoization strategy which caches and reuses geometric data for each surface, such as surface points and tangent vectors at quadrature nodes (detailed in Section 3.4). The second term is a correction term that accounts for any jump discontinuities in the GWN field over the surface.
@@
  • Far-field: For points sufficiently far from the surface, i.e. outside a bounding box encompassing the surface, the number of intersections is known to be zero. As such, we evaluate the GWN directly with our unadjusted boundary formulation.
@@
The specific form of Equation 6 assumes that the correction term is determined by the $z$-axis. However, the intersection test can be performed with an arbitrary oriented line by subsequently rotating the surface so that the positive direction of the line is parallel to the positive $z$-axis. As intersections with lines that are near-tangent to the surface are the most difficult to identify and resolve, we heuristically select a line with a direction equal to an average surface normal, which we compute for each surface as preprocessing. In cases where the symmetry of the surface results in an average normal with near-zero magnitude, we instead select the line randomly.
@@ 3.2. Reformulation with Stokes’ Theorem
The connection between the 3D solid angle and an associated line integral formulation has been independently explored in several application contexts, including physical optics (Asvestas, 1985), soft matter physics and electrostatics (Binysh and Alexander, 2018), and perhaps most directly in differential geometry (Borodzik et al., 2017). For example, there are many contexts in which the solution to specific line integrals over closed 3D contours can be more efficiently evaluated by means of computing the (often approximated) solid angle of some interior surface (Binysh and Alexander, 2018). Common among these works is either the selection or presupposition of a particular line $L$ which specifies the particular form of the boundary formulation, as is the case in the presented work. Indeed, the work of Asvestas (1985) presents a solution strategy which most closely mirrors our own, where signed intersections between the line $L$ and a planar surface are distinguished according to the location of the intersection, with the numerical result further adjusted through an analytically derived correction term.
Directly applying such methods to the present graphics context is restricted by the fact that the targeted geometry is more complex and cannot be a priori discretized (see Section 2.2). However, we have the advantage of being able to select the line $L$ which is most computationally advantageous, and further manipulate the boundary curves of the surface to ensure our algorithm’s robustness. As such, we present our own derivation of these principles in differential geometry from the ground up, as doing so leads to our proposed computational treatment of general trimmed NURBS surfaces.
@@
(8) $\nabla\times F=\frac{1}{4\pi}\cdot\frac{\vec{x}}{\lVert\vec{x}\rVert^{3}}.$
Recall that we have made the assumption that the query point $q$ is located at the origin, and so this integrand is continuous on the surface whenever $\vec{x}\neq 0$, i.e. the query point does not coincide with the surface (we treat this case explicitly in Section 3.3).
There are many distinct vector fields $F$ that satisfy Equation 8, and so one could evaluate the GWN for an arbitrary surface with the following 1D line integral,
@@
assuming that the theoretical conditions of Stokes’ theorem are met for the surface $S$.
This reformulation of the problem statement comes with a number of computational advantages. First, evaluating such an integral via numerical quadrature along the lower-dimensional space naturally improves the efficiency of the total method. This is a well-established advantage of Stokes’ theorem, and is the basis of a method proposed by Gunderman et al. (2021) for evaluating arbitrary integrals on trimmed NURBS surfaces. In this problem context, we can evaluate the “antiderivative” $F$ exactly for the relevant $\nabla\times F$, further reducing the number of quadrature nodes in a scheme of fixed order $n$ from $O(n^{2})$ to $O(n)$.
Second, such a boundary reformulation fundamentally improves the stability of the ensuing quadrature scheme, as query points which are near to the surface but not near the boundary are not subject to the same near-singular behavior in the corresponding integrand. The resulting line integrals are additionally more amenable to a geometrically adaptive quadrature method, the computational burden of which is dramatically reduced by considering subdivision of 3D space curves instead of entire trimmed surfaces.
However, not every surface can be treated directly by the formulation in Equation 9. To meet the conditions of Stokes’ theorem, both the original vector field and its antiderivative must be continuous—not just on its domain of integration (the boundary of the surface) but also on the original domain of integration (the surface itself). Notably, the relevant vector field $F$ is discontinuous whenever $x^{2}+y^{2}=0$ on the surface, i.e. whenever the surface intersects a line extending in both $z$ directions from the query point. In such cases, we cannot mathematically justify the direct use of Stokes’ theorem, even if the resulting line integral is itself continuous.
We note that the specific antiderivative in Equation 9, characterized by a line of discontinuities on the $z$-axis, was chosen arbitrarily. Indeed, one can analogously define a family of antiderivatives $F$, each characterized by a different line of discontinuities.
@@
Example configurations for surface, query point, and line of discontinuity.
In Figure 5, we see examples of different configurations of the surface and the line. As shown in Figure 5 (b), it is practical to view different choices for $F$ not as explicit vector fields, but as different rotations of the NURBS patch via its control points so that the desired line of discontinuities is aligned to the $z$-axis. In this way, applying Stokes’ theorem requires choosing a line which does not intersect the surface, and rotating the surface accordingly. In practice, one cannot always identify such a line for query points which are nearer to the surface, nor does one necessarily exist (see Figure 5 (c)).
@@ 3.2.1. Far-field: Direct Application of Stokes’ Theorem
While our implementation first checks an axis-aligned bounding box to identify far-field scenarios, other convex shapes containing the shifted surface can also be used to guarantee a non-intersecting line $L$, again followed by the appropriate rotation of the surface. For example, our implementation also checks a precomputed oriented bounding box for each NURBS surface, allowing more points to be classified as far-field. In any case, these are the most commonly evaluated forms of the integral, as an arbitrary query point will be considered far from all but a few surfaces in a typical CAD model.
The validity of this far-field formulation can be made more intuitive by making concrete a property of the GWN introduced in Section 2.3 – namely, that the GWN field of a surface can, for far away points, be defined entirely by its boundary. More specifically, any two surfaces $S$ and $S^{\prime}$ which share a boundary $\partial S=\partial S^{\prime}$ generate an identical GWN field at all points outside a volume implied by $S\cup S^{\prime}$. Furthermore, if one were to continuously deform $S$ into $S^{\prime}$, then the value of the GWN at a fixed query point $q$ inside this volume would remain constant up until the surface passes through $q$, at which point the GWN value would jump by one. We depict this scenario explicitly in Figure 6, which further illustrates the specific form of Equation 6. Although the GWN field is determined almost entirely by boundary data (and completely so for far-away points), we must still account for the influence of the internal surface; as we will show, it affects the GWN value only by a constant.
Query Left Right
Point Shape GWN Shape GWN Difference
$q_{1}$ -0.20139 -0.20139 0.0
$q_{2}$ 0.62912 -0.37088 1.0
$q_{3}$ 0.27805 0.27805 0.0
Figure 6. We illustrate a property of the GWN for two surfaces which share a boundary, viewed from different angles. In general, any two surfaces which share a boundary generate the same GWN field, modulo some integer which depends on the orientation of the query point relative to the surface.
Illustrative example of 3D GWN algorithm The GWN for 2 surfaces which share a boundary is shown on a 2D slice. The two surfaces generate an identical GWN field except in the region between them.
@@ 3.2.2. Near-field: Analytically Adjusted Stokes’ Theorem
If the surface bounding box does contain the origin (i.e. the query point is inside the surface bounding box), we can no longer guarantee that there exists a line $L$ which does not intersect the surface. Instead, we must explicitly test for signed intersection between some oriented line $L$ and the surface. We consider cases to be “near-field” when all intersections between $L$ and the surface occur away from the surface edges and other degeneracies. When these conditions are not met, the query point is handled as an “edge-case”.
@@
Naturally, distinguishing the near-field and edge cases requires first identifying the location of all intersections between the surface and the line in the parameter space of the surface. As noted in Section 2, there are many numerical ray casting algorithms one could apply to find signed intersections between this line and the surface, and in principle, any can be used for the purposes of evaluating the GWN. We describe our own geometric subdivision approach in Section 5.1.1, which in contrast to many algebraic alternatives, takes advantage of the broader GWN algorithm’s high tolerance for imprecision in the intersection routine.
@@
Because any intersection between $L$ and the surface violates the conditions of Stokes’ theorem, direct application of the boundary formulation is theoretically unfounded and, in general, evaluates to an incorrect GWN. Our proposed correction term in Equation 6 permits evaluation of the GWN using the same boundary integral formulation as Equation 9, provided that for each intersection (if any), the value $0.5$ is added or subtracted according to the orientation of the surface and line at the point of intersection to compensate for the jump discontinuity present in the GWN field along the surface.
@@
This is illustrated in Figure 7 (bottom), showing that the difference between the ground truth GWN and the specific boundary formulation in Equation 9 is always, perhaps unexpectedly, half-integer valued. As a result, the true GWN can always be recovered from the boundary formulation by identifying the proper half-integer value.
Figure 7. We illustrate a property of the GWN in both 2D (top) and 3D (bottom). The difference between the GWN for a curved surface (left) and that of a particular boundary term (middle) is always a constant value. In 2D, when this boundary term is the GWN of the curve’s linear closure, the difference is integer valued. In 3D, when this boundary term is the integral component of Equation 6, the difference is always half-integer valued. This suggests a method for GWN evaluation by evaluating the boundary term, and adjusting it by identifying the proper constant.
Illustrative example of boundary formulation A 2D and 3D example of the GWN for a curved shape, the scalar field generated by the associated boundary formulation, and the difference between them. In both cases, the difference is a piecewise-constant field.
@@
In some sense, the existence of an analytic fix that depends only on the orientation of the query point relative to the surface is a natural consequence of the double-layer potential which, when integrated along the surface, defines the GWN scalar field (Barill et al., 2018; Bang et al., 2023). The use of this specific correction term is made more intuitive by means of analogy in two spatial dimensions (with a direct proof provided in Appendix A). As summarized in Section 2.3 and demonstrated in Figure 7, the GWN for an arbitrary curve in 2D (left) can always be computed as the difference between an integer winding number for the curve closed by a straight line (right) and a generalized winding number defined by that straight line closure (middle). Crucially, the GWN of the closure is uniquely defined by the two boundary points of the original curve. From the lens of our 3D GWN framework, the GWN of the closure can be thought of as a “boundary term” which must be “corrected” by the integer winding number of the closed shape.
@@
To make the comparison explicit, we see that in both 2D and 3D, the value of the GWN is the sum of a fractional component which does not depend on any interior geometry, and an integer component which does. Although the fractional component in 2D reduces to a subtended angle, while the 3D equivalent is not the GWN of any particular shape, both integer components can be thought of as determined by the result of a line intersection test. Most importantly, points which are known to be far from the shape do not require any correction at all. Returning our focus to 3D, we emphasize that although our line-surface intersection test does come at additional cost, it is necessary only for the surfaces nearest to the query point.
@@ 3.2.3. Edge Cases: Disk Extraction via Trimming Curves
Edge-Case Evaluation for 3D GWN Several examples of edge-case handling are shown, including the line of discontinuities intersecting (a) the boundary of the surface, (2) a cusp on the surface, (c) at a point of tangency with the surface. In the first two cases, a parameter-space disk is removed from the surface and processed separately. In the third, the surface is randomly rotated and processed again.
We now consider the remaining edge cases for which the query point is inside the surface bounding box, but the correction term in Equation 6 cannot be applied. As shown in Figure 8, this occurs when the line of discontinuities intersects the surface at a boundary, at a cusp, or is tangent to the surface. We treat all edge cases by either finding a new line (in the case of tangencies) or by removing from the surface a disk of radius $r$ in parameter space around the point of intersection. This guarantees that the line does not intersect the surface at the otherwise problematic intersection point, and permits evaluation via Equation 6. The extracted disk is then fed back into the original algorithm for reprocessing. Since the produced disk has a much smaller bounding box than the original surface, the computational cost of the reprocessing is significantly reduced.
@@
The only edge case which is not well-handled by this strategy is a line that is tangent (or near-tangent) to the surface, as the surface remaining after the disk is removed may still intersect the line. In these cases, we instead rotate the surface (i.e. consider a different line) and feed the rotated surface into the original algorithm (see Figure 8 (c)). While this is the most computationally expensive type of edge case, it is also the least common due to our choice of the cast ray to be an approximate surface normal (See Section 5.1.1).
@@
3.3. GWN for Points Coincident with the Surface
@@ 3.4. Accelerated Adaptive Quadrature
In our implementation, we utilize a recursive subdivision approach, applying a fixed-order Gauss-Legendre quadrature rule at each level of recursion. We stop the recursion when the difference between the evaluated integral on the entire curve and the sum of the integrals on the two halves is less than a user-defined quadrature tolerance $\epsilon_{q}$. Thus, in most cases, $\epsilon_{q}$ determines how many digits of precision are achieved in the computed GWN field, although it is possible for more lenient values of $\epsilon_{q}$ that the two terms will agree at an incorrect value. Using a Gaussian rule at each recursion level is particularly useful as it is capable of maintaining high precision even without specific structure in the integrand. We specifically chose this recursive subdivision strategy because it effectively concentrates quadrature nodes near unstable regions of the integrand, i.e. when the query point is close to the 3D space curve which defines the domain of integration. Elsewhere, where the integrand is stable, we maintain the typically high precision of Gaussian quadrature without additional computational expense.